Set K contains 50 elements, set J contains 66 elements, and 14 elements are common to both sets. Find n (K U J).
=50+66-14
=102

Did I do this problem correctly? If not, can someone help me do it correctly? I appreciate it!

Answer by solver91311(5072)



Yes, you did it correctly in that you obtained the correct answer. However, there is another way to look at it that might make more sense and will help you later with more complex overlapping set problems in the future.

Draw two overlapping circles. One of the circles represents set K and the other represents set J, and the overlap region in the middle represents the elements that the two sets have in common.

Write the number 14 in the overlap region. Then the number of elements of K that go in the other part of the K circle is 50 - 14 = 36, so write 36 in the region of the K circle that doesn't overlap the J circle. Likewise, the number of elements of J that are not in the overlap is 66 - 14 = 52 -- write that in the non-overlap region of the J circle.

Now, the number of elements in the union of the two sets is the sum of the three numbers you wrote: 36 + 52 + 14 = 102.


John